# Runge–Kutta Physics Informed Neural Networks: Formulation and Analysis

### v1

@@ 4.3.2 Discontinuous Galerkin methods
Following [MakrN2006posteriori], we define the reconstruction $\widehat{U}$ of the dG approximation $U,$ the analogue of the collocation approximation, by extended interpolation at the Radau nodes $t_{ni}=t_{n}+c_{i}k,0<c_{1}<\dotsb<c_{q}=1,$
@@ 4.4.2 Maximal regularity of Lobatto IIIA methods
Now, the integrand $\pi:=\big{(}\widehat{U}(\cdot)-I_{q-1}\widehat{U}(\cdot),v^{\prime}(\cdot))$ in (98) is a polynomial of degree at most $2q-3;$ therefore, $\pi$ is integrated exactly by the Lobatto quadrature formula with $q$ nodes. Furthermore, $\pi$ vanishes at the quadrature nodes $t_{n1},\dotsc,t_{nq}$. Thus, (98), and thus also (97), is indeed valid.
@@ 5.1 Problem Setup
with additional boundary conditions for $t\in[0,T].$ In the numerical experiments below the boundary conditions are either Neumann (for the heat equation) or Dirichlet (for the wave equation). The formulation of the methods can be directly extended to nonlinear evolution equations with the obvious modifications in the loss functionals, see [AMN2].
@@ 5.2 Collocation Runge–Kutta Formulation
Furthermore, we let $10=\tilde{c}_{0}<\tilde{c}_{1}<\dots<\tilde{c}_{q}=$ 1 be auxiliary points as introduced in Section 2.4.2, and
@@
$I_{q-1}L\hat{u}(x,t)=\sum_{j=1}^{q}\ell_{j}\left(\frac{t-t_{n}}{k_{n}}\right)\sum_{i=0}^{q}\tilde{\ell}_{i}\left(\frac{t_{nj}-t_{n}}{k_{n}}\right)Lu(x,\tilde{t}_{ni})$ (115)
@@
$\begin{split}\zeta(x,t_{nj})=&\hat{u}_{t}(x,t_{nj})+I_{q-1}L\hat{u}(x,t_{nj})\\ =&\sum_{i=0}^{q}\big{(}k_{n}^{-1}\tilde{\ell}_{i}^{\prime}(c_{j})u_{m}(x,\tilde{t}_{nj})-\tilde{\ell}_{i}(c_{j})Lu_{m}(x,\tilde{t}_{nj})\big{)},\quad x\in\varOmega,\end{split}$ (118)
@@ 5.3 Artificial Neural Network Representation
We approximate the solution of our problem within the neural network function space. The objective is to find $\theta^{\star}\in\Theta$ such that the $m$-th output $U_{m}$ of the neural network approximates the target function $u(\cdot,\tilde{t}_{ni})$. Specifically, we require that the neural network output satisfies the following approximation.
@@
$\mathcal{C}_{\varOmega}[\theta]=\frac{\mathrm{Vol}(\varOmega)}{R}\sum_{m=1}^{M}\sum_{r=1}^{R}\sum_{n=1}^{N}\int_{J_{n}}\zeta(x_{r},t)^{2}\,{\mathrm{d}}t,$ (121)
where $\zeta$ is given by (119) and the set of points $\{x_{r}\}_{r=1}^{R}$ is generated using Sobol’s low-discrepancy sequences, see e.g., [Caflisch_review_1998]. This sampling approach is more efficient for achieving a uniform coverage of the space in higher–dimensonal settings. The integral in time is be computed exactly by applying integration rules which are exact for polynomials of degree $22q-$ 2 . In the case where we have Gauss or Radau collocation points this integral is just
@@
where $x_{r^{\prime}}$ are Sobol points on the boundary ${\partial\varOmega_{s}}.$ In case we use other than Dirichlet boundary conditions, this term is modified accordingly. The deep learning approximation of the problem is characterized by the minimization of the sum of the cost functionals with respect to model parameters $\theta$
@@ 5.4 Applications
Notice that, in the case of Radau and Lobatto methods, the points $10=\tilde{c}_{0}<\tilde{c}_{1}<\dotsb<\tilde{c}_{3}=$ 1 include the collocation points $c_{1},\dotsc,c_{3}$. To ensure a fair comparison, we use four times the value of $n$ for the uniform sampling case compared to the Runge–Kutta collocation schemes with $q=3$.
@@ 5.4.2 Heat Equation with a Discontinuous Initial Value
We replace the initial condition from the previous test case with a discontinuous one. Specifically, we consider
@@
The heat equation exhibits a notable smoothing property, which is not always maintained by various time discretization methods, see, e.g., [Luskin_Ran_CN_1982], [Thomee_book_2006]. The smoothing behavior of time discretizations is a quite subtle topic, as it can influence the performance of methods when both diffusion and transport phenomena are present, particularly in the context of nonlinear PDEs and Navier-Stokes equations, see [Glow_theta1987]. Although smoothing eventually occurs primarily due to the diffusion induced by the Quasi-Monte Carlo method used for spatial discretization, it is worth noting, as shown in Fig. 5, that the behavior of various discretization methods aligns with the known predictions of the corresponding Runge-Kutta schemes. In Fig. 4, we compare the various schemes we compare the various schemes with respect to heat conservation.
@@
Figure 5: Smoothing effect for the heat equation. As expected, Gauss and Lobatto methods have oscillating behaviour close to initial times for discontinuous data. The full smoothing effect of Radau methods is evident.
@@ 5.4.3 Wave Equation
$\begin{pmatrix}u\\ v\end{pmatrix}_{t}+\begin{pmatrix}0&-\varDelta\\ -I&0\end{pmatrix}\begin{pmatrix}u\\ v\end{pmatrix}=\begin{pmatrix}0\\ 0\end{pmatrix},$ (134)

### v2

@@ 4.3.2 Discontinuous Galerkin methods
Following [MakrN2006posteriori], we define the reconstruction $\widehat{U}$ of the dG approximation $U,$ the analogue of the collocation approximation, by extended interpolation at the Radau nodes $t_{ni}=t_{n}+c_{i}k_{n},0<c_{1}<\dotsb<c_{q}=1,$
@@ 4.4.2 Maximal regularity of Lobatto IIIA methods
Now, the integrand $\pi:=\big{(}\widehat{U}(\cdot)-I_{q-1}\widehat{U}(\cdot),v^{\prime}(\cdot))$ in (98) is a polynomial of degree at most $2q-3;$ therefore, $\pi$ is integrated exactly by the Lobatto quadrature formula with $q$ nodes. Furthermore, $\pi$ vanishes at the quadrature nodes $t_{n1},\dotsc,t_{nq}$. Thus, (98), and hence also (97), is indeed valid.
@@ 5.1 Problem Setup
with additional boundary conditions for $t\in[0,T].$ In the numerical experiments below the boundary conditions are either Neumann (for the heat equation) or Dirichlet (for the wave equation). The formulation of the methods can be directly extended to nonlinear evolution equations with the obvious modifications in the loss functionals; see [AMN2].
@@ 5.2 Collocation Runge–Kutta Formulation
Furthermore, we let $10=\tilde{c}_{0}<\tilde{c}_{1}<\dotsb<\tilde{c}_{q}=$ 1 be auxiliary points as introduced in Section 2.4.2, and
@@
$I_{q-1}L\hat{u}(x,t)=\sum_{j=1}^{q}\ell_{j}\left(\frac{t-t_{n}}{k_{n}}\right)\sum_{i=0}^{q}\tilde{\ell}_{i}\left(\frac{t_{nj}-t_{n}}{k_{n}}\right)Lu(x,\tilde{t}_{ni}).$ (115)
@@
$\begin{split}\zeta(x,t_{nj})=&\hat{u}_{t}(x,t_{nj})+I_{q-1}L\hat{u}(x,t_{nj})\\ =&\sum_{i=0}^{q}\big{(}k_{n}^{-1}\tilde{\ell}_{i}^{\prime}(c_{j})u(x,\tilde{t}_{nj})-\tilde{\ell}_{i}(c_{j})Lu(x,\tilde{t}_{nj})\big{)},\quad x\in\varOmega,\end{split}$ (118)
@@ 5.3 Artificial Neural Network Representation
We approximate the solution of our problem within the neural network function space. The objective is to find $\theta^{\star}\in\Theta$ such that the $m$-th output $U_{m}$ of the neural network approximates the target function $u_{m}(\cdot,\tilde{t}_{ni})$. Specifically, we require that the neural network output satisfies the following approximation
@@
$\mathcal{C}_{\varOmega}[\theta]=\frac{\mathrm{Vol}(\varOmega)}{R}\sum_{m=1}^{M}\sum_{r=1}^{R}\sum_{n=1}^{N}\int_{J_{n}}\zeta_{m}(x_{r},t)^{2}\,{\mathrm{d}}t,$ (121)
where $\zeta_{m}$ denotes the $m$-th component of $\zeta$ which is given by (119) and the set of points $\{x_{r}\}_{r=1}^{R}$ is generated using Sobol’s low-discrepancy sequences; see, e.g., [Caflisch_review_1998]. This sampling approach is more efficient for achieving a uniform coverage of the space in higher–dimensonal settings. The integral in time is computed exactly by applying integration rules which are exact for polynomials of degree $22q-$ 2 . In the case where we have Gauss or Radau collocation points this integral is just
@@
where $x_{r^{\prime}}$ are Sobol points on the boundary ${\partial\varOmega_{s}}.$ In case we use other than Dirichlet boundary conditions, this term is modified accordingly. The deep learning approximation of the problem is characterized by the minimization of the sum of the cost functionals with respect to model parameters $\theta$,
@@ 5.4 Applications
Notice that, in the case of Radau and Lobatto methods, the points $10=\tilde{c}_{0}<\tilde{c}_{1}<\tilde{c}_{2}<\tilde{c}_{3}=$ 1 include the collocation points $c_{1},\dotsc,c_{3}$. To ensure a fair comparison, we use four times the value of $n$ for the uniform sampling case compared to the Runge–Kutta collocation schemes with $q=3$.
@@ 5.4.2 Heat Equation with a Discontinuous Initial Value
We replace the initial value from the previous test case with a discontinuous one. Specifically, we consider
@@
The heat equation exhibits a notable smoothing property, which is not always maintained by various time discretization methods; see, e.g., [Luskin_Ran_CN_1982], [Thomee_book_2006]. The smoothing behavior of time discretizations is a quite subtle topic, as it can influence the performance of methods when both diffusion and transport phenomena are present, particularly in the context of nonlinear PDEs and Navier–Stokes equations; see [Glow_theta1987]. Although smoothing eventually occurs primarily due to the diffusion induced by the Quasi-Monte Carlo method used for spatial discretization, it is worth noting, as shown in Fig. 5, that the behavior of various discretization methods aligns with the known predictions of the corresponding Runge–Kutta schemes. In Fig. 4, we compare the various schemes with respect to heat conservation.
@@
Figure 5: Smoothing effect for the heat equation. As expected, Gauss and Lobatto methods have oscillating behavior close to initial times for discontinuous data. The full smoothing effect of Radau methods is evident.
@@ 5.4.3 Wave Equation
$\begin{pmatrix}u\\ v\end{pmatrix}_{t}+\begin{pmatrix}0&-I\\ -c^{2}\varDelta&0\end{pmatrix}\begin{pmatrix}u\\ v\end{pmatrix}=\begin{pmatrix}0\\ 0\end{pmatrix},$ (134)